### v1

@@ 6.1.1 Overview
Given a cluster $S\subseteq V$, we would like to check if its boundary edges $B_{S}=E(S,V\setminus S)$ expands within G[S]G[S]. [^17] The formal argument is performed in the subdivision graph (i.e.proving that $X_{B_{S}}$ expands in $G^{\prime}[S^{\prime}]$). See Section 6.1.2 for details. For this overview, we omit these details for simplicity. If they are not, we would like to find a sparse cut $(A,S\setminus A)$ with respect to the corresponding weighting $\mu_{S}=\mathbbm{1}_{B_{S}}$, that certifies that the boundary edges do not expand. We will then recur on both $A$ and $S\setminus A$ (with the new weightings $\mu_{A}=\mathbbm{1}_{B_{A}}$ and similarly $\mu_{S\setminus A}=\mathbbm{1}_{B_{S\setminus A}}$). We stop when we find a sub-cluster in which the boundary edges expand (equivalently, the sub-cluster expands with respect to the weighting of the boundary edges). This can be facilitated using Theorem 3.9. Note that when we recur into a sub-cluster (e.g., $A\subseteq S$), the new cut edges $E(S,A\setminus S)$ are now considered boundary edges of the sub-clusters $A$ and $S\setminus A$. However, because the cut was sparse, the capacity of these cut edges is small compared to the already-present boundary edges, which ensures that the total capacity of the boundary edges decreases. Note that we must carefully adjust the target expansion $\phi$ when we invoke Theorem 3.9 on the different sub-clusters, to ensure that the final sub-clusters $S_{i}\subseteq S$, $\Omega\!\left(\frac{1}{f(S_{i})\cdot\log n}\right)$-expand with respect to the boundary edges.
@@
We can imagine the sequence of cuts and recursions as a binary tree whose root is $S$, and where in each step, the children of the current node are the two sides of the cut. Because the cut is balanced, we can ensure that the depth of the tree is poly-logarithmic, so the total running time is $\tilde{O}(m)$.
@@
As for the second result of Theorem 6.2, we would like to use the flows that are promised by Remark 3.10 to be able to route from the cut edges $E(A,S\setminus A)$ in each iteration, to the boundary edges of $A$ or of $S\setminus A$. In this manner, we could route the flow from the inter-cluster edges $F$ to the boundary edges $B_{S}$ by going over the binary tree from the leaves upwards: each time we encounter a cut $(A,D\setminus A)$ of the current sub-cluster $D\subseteq S$, we route the mass currently residing on the cut edges to the boundary edges of either $A$ or $D\setminus A$. Note that the demands on the boundary edges accumulate as we perform this step. In particular, after accounting for all relevant details, we obtain that routing a unit of flow from each edge of $E(A,D\setminus A)$ to $A$ adds $O\left(\frac{1}{f(D)}\right)$ units of load to the boundary edges of $A$ (and similarly if we route to the boundary of $D\setminus A$). Assuming the load on the boundary edges of $A$ and on the cut edges $E(A,D\setminus A)$ (accumulated in the lower levels of the binary tree) was $\alpha$, this means that when we go upwards in the tree, the load on the boundary of $D$ will now be $\alpha^{\prime}=\alpha\cdot\left(1+\frac{O(1)}{f(D)}\right)$.
@@
If we could always route to the side with the smaller number of vertices (i.e., route to $A$ if $|A|\leq|D\setminus A|$ and to $|D\setminus A|$ otherwise), this approach would work. The reasoning is as follows. Since $f(D)$ is defined as $c_{f}\cdot\log\left(\frac{\sigma}{|D|}\right)\log\log n$ and $\sigma\geq\frac{3}{2}|S|$, it follows that if $|D|\leq|S|\cdot 2^{-j}$ for some $j\geq 1$, it implies $f(D)=\Omega(j\log\log n)$. This ensures that for each edge $e$, if we consider the sequence of clusters $D_{i}$ (along the binary tree) in which the load on $e$ increases (i.e., when $e$ is a boundary edge on the side with fewer vertices), the product $\prod_{i}{\left(1+\frac{1}{f(D_{i})}\right)}$ behaves like $\prod_{j=1}^{\log n}{\left(1+\frac{1}{j\log\log n}\right)}$, so it remains bounded by a constant. Hence, the total load accumulated on each edge, determined by this product, is constant.
However, there is a problem with this approach. Specifically, as detailed in Remark 3.10, due to the internal algorithms of Theorem 3.9, the returned sparse cut $(A,S\setminus A)$ of a cluster $S$ is created from a sequence of cuts $\{S_{t}\}_{t=0}^{T^{\prime}-1}$, where if we define $A_{0}=S$, $A_{t+1}=A_{t}\setminus S_{t}$, we have that $S_{t}\subseteq A_{t}$ is the cut we remove from the leftover cluster at step $t$. All of these cuts are sparse (i.e., $(S_{t},A_{t}\setminus S_{t})$ is a sparse cut of $A_{t}$), and additionally we can route from the cut edges to both sides ($S_{t}$ and $A_{t}\setminus S_{t}$).
@@
Ultimately, when $A=S\setminus\left(\bigcup_{t}{S_{t}}\right)$, we can route flow from the cut edges $(A,S\setminus A)$ to the side $S\setminus A$ (see Figure 6(2a)(2a)). However, if we attempt to route instead towards $A$, we may incur congestion on the same edges multiple times (specifically, $T^{\prime}$ times), preventing a routing with sufficiently low congestion. We emphasize that this is not an issue in $S\setminus A$, because each cut is removed from the graph after it is processed: we route from $E(S_{t},A_{t}\setminus S_{t})$ to $S_{t}$ within $G[S_{t}]$, and then $S_{t}$ is removed from the remaining graph. As a result, the corresponding flow routings remain disjoint (see Figure 6(2a)(2a)).
@@
This routing is good for us if $|S\setminus A|\leq|A|$, but we have no such guarantee from Theorem 3.9 (as the theorem only bounds $\mu(S\setminus A)$, which is the number of boundary edges). Instead, we choose to do something slightly more nuanced: we stop the sequence of cuts $\{S_{t}\}_{t=0}^{T^{\prime}}$ early at some $T_{0}<T^{\prime}$, as soon as its size $\left|\bigcup_{t=0}^{T_{0}}{S_{t}}\right|$ is a constant fraction of $|S|$. In other words, we find the first $T_{0}$ such that $\left|\bigcup_{t=0}^{T_{0}}{S_{t}}\right|=\Theta(|S|)$. Denote $A=S\setminus\left(\bigcup_{t=0}^{T_{0}}{S_{t}}\right)$. Assuming that at this step, both sides of the cut are of size $\Theta(|S|)$, this is enough to ensure that the recursion depth is poly-logarithmic (even though $\mu(S\setminus A)=\mu\left(\bigcup_{t=0}^{T_{0}}{S_{t}}\right)$ may be arbitrarily small), and allows us to bound $|S\setminus A|=\left|\bigcup_{t=0}^{T_{0}}{S_{t}}\right|=\Theta(|S|)$ (see Figure 6(2b)(2b)).
@@
The final obstacle with this approach is that it may be that while $\left|\bigcup_{t=0}^{T_{0}-1}{S_{t}}\right|$ is small (say $o(|S|)$), it’s possible $\left|\bigcup_{t=0}^{T_{0}}{S_{t}}\right|$ is already too large (say $(1-o(1))\cdot|S|$). In this case we perform two steps: we denote $A_{1}\mathrel{\mathop{:}}=S\setminus\left(\bigcup_{t=0}^{T_{0}-1}{S_{t}}\right)$, which is a cut of $S$, and then further cut $A_{1}$ into $(A_{2}\mathrel{\mathop{:}}=A_{1}\setminus S_{T_{0}},S_{T_{0}})$. Routing from $E(A_{1},S\setminus A_{1})$ to boundary of the side $S\setminus A_{1}$ is done like before (as mentioned, routing to the side $S\setminus A_{1}$ does not incur multiple congestion on the same edges, because in $S\setminus A_{1}=\bigcup_{t=0}^{T_{0}-1}{S_{t}}$, the cuts are removed at each step). Routing from $E(A_{2},S\setminus A_{2})$ to the boundary of $A_{2}$ is possible because $S_{T_{0}}$ now represents a single inner cut rather than a sequence of cuts. Consequently, we do not incur the congestion multiple times (see Figure 6(2c)(2c)).
@@
In total, this partitioning scheme guarantees that we always route from the cut edges to the boundary edges of a side whose size is at most $c\cdot|S|$. This suffices to show that the total accumulated load on each boundary edge remains bounded by a constant.
@@
We are going to apply Theorem 3.9 on the cluster $S\subseteq V$. However, as discussed in Section 6.1.1, we require a slightly refined corollary of this theorem. Following the intuition presented in Section 6.1.1, one should think about the measure $\nu$ used in the corollary below as simply being the number of vertices $\nu(A)=|A|$. [^18] As will be described below, we use Corollary 6.5 on the subdivision graph and $\nu(A)$ will actually be the number of “non-split vertices” in $A$, i.e., $\nu(A)=|A\cap V|$. Recall the notion of routing from a cut to a set from Section 3.3.
@@
Given a graph $G=(V,E)$ of $m$ edges, a parameter $\phi>0$, and two weightings $\mu,\nu:V\to\mathop{\mathbb{R}}_{\geq 0}$ of the vertices, there exists a randomized algorithm which takes $\tilde{O}(m)$ time and must end in one of the following cases:
@@
    We find a cut $(A,V\setminus A)$ of $V$ with $\mu$-expansion $\Phi^{\mu}_{G}(A,V\setminus A)=O(\phi\log n)$, and $\mu(V\setminus A)=\Omega\left(\frac{\mu(V)}{\log n}\right)$. Additionally $\nu(V\setminus A)\leq\frac{1}{4}\nu(V)$. There exists a flow routable with congestion $O(1)$ in $G[V\setminus A]$, such that each cut edge of $E(A,V\setminus A)$ is a source of one unit of flow [^19] Formally, there is a flow in $G[V\setminus A]$ routable with congestion $O(1)$, such that each $v\in V\setminus A$ sends $\deg_{G}(v)-\deg_{G[V\setminus A]}(v)$ units of flow and receives at most $O(\phi\log n\cdot\mu(v))$ units. The rest of the flows in this corollary are defined similarly. and each $v\in V\setminus A$ receives at most $O(\phi\log n\cdot\mu(v))$ units.
@@
    We find a cut $(A,V\setminus A)$ of $V$ with $\mu$-expansion $\Phi^{\mu}_{G}(A,V\setminus A)=O(\phi\log n)$. Additionally $\frac{1}{4}\nu(V)\leq\nu(V\setminus A)\leq\frac{1}{2}\nu(V)$. There exists a flow routable with congestion $O(1)$ in $G[V\setminus A]$, such that each cut edge of $E(A,V\setminus A)$ is a source of one unit of flow and each $v\in V\setminus A$ receives at most $O(\phi\log n\cdot\mu(v))$ units.
@@
    We find a cut $(A_{1},V\setminus A_{1})$ of $V$ and a cut $(A_{2},A_{1}\setminus A_{2})$ of $A_{1}$, with $\mu$-expansion $\Phi^{\mu}_{G}(A_{1},V\setminus A_{1}),\Phi^{\mu}_{G}(A_{2},A_{1}\setminus A_{2})=O(\phi\log n)$. Additionally, $\mu(A_{2})\geq\frac{1}{4}\mu(V)$ and $\nu(A_{1}\setminus A_{2})\geq\frac{1}{4}\nu(V)$. There exists a flow routable with congestion $O(1)$ in $G[V\setminus A_{1}]$, such that each cut edge of $E(A_{1},V\setminus A_{1})$ is a source of one unit of flow and each $v\in V\setminus A_{1}$ receives at most $O(\phi\log n\cdot\mu(v))$ units. There also exists a flow routable with congestion $O(1)$ in $G[A_{2}]$, such that each cut edge of $E(A_{2},A_{1}\setminus A_{2})$ is a source of one unit of flow and each $v\in A_{2}$ receives at most $O(\phi\log n\cdot\mu(v))$ units.
@@
    We find a cut $(A,V\setminus A)$ with $\mu$-expansion $\Phi^{\mu}_{G}(A,V\setminus A)=O(\phi\log n)$. Additionally, $0<\mu(V\setminus A)\leq\frac{\mu(V)}{2}$ and $\nu(V\setminus A)\geq\frac{1}{2}\nu(V)$. There exists a flow routable with congestion $O(1)$ in $G[A]$, such that each cut edge of $E(A,V\setminus A)$ is a source of one unit of flow and each $v\in A$ receives at most $O(\phi\cdot\mu(v))$ units.
@@
    We find a cut $(A,V\setminus A)$ with $\mu$-expansion $\Phi^{\mu}_{G}(A,V\setminus A)=O(\phi\log n)$, and with high probability, $G[A]$ is an $\Omega(\phi)$-expander with respect to $\mu$. Additionally, $\nu(V\setminus A)\leq\frac{1}{2}\nu(V)$. There exists a flow routable with congestion $O(1)$ in $G[V\setminus A]$, such that each cut edge of $E(A,V\setminus A)$ is a source of one unit of flow and each $v\in V\setminus A$ receives at most $O(\phi\log n\cdot\mu(v))$ units. Additionally, there exists a flow routable with congestion $O(1)$ in $G[A]$, such that each cut edge of $E(A,V\setminus A)$ is a source of one unit of flow and each $v\in A$ receives at most $O(\phi\cdot\mu(v))$ units.
@@
We apply Theorem 3.9 on $G$. If it terminated with case (1), we are done. Otherwise, first assume it terminated with case (2). Recall from Remark 3.10 that $\bar{A}$ is given as $\bar{A}=\bigcup_{t}{S_{t}}$, where if we look at the sequence of vertex sets $A_{0}=V,A_{t+1}=A_{t}\setminus S_{t}$, each $(S_{t},A_{t+1})$ is a sparse cut. We split into cases according to $\nu(\bar{A})$. If $\nu(\bar{A})\leq\frac{1}{4}\nu(V)$, we terminate with case (2a) (see Figure 6).
Next, assume that $\nu(\bar{A})>\frac{1}{4}\nu(V)$. We consider the first $T_{0}\geq 0$ such that $\nu\left(\bigcup_{t=0}^{T_{0}}{S_{t}}\right)>\frac{1}{4}\nu(V)$ (i.e., $\nu\left(\bigcup_{t=0}^{T_{0}-1}{S_{t}}\right)\leq\frac{1}{4}\nu(V)$). If $\nu\left(\bigcup_{t=0}^{T_{0}}{S_{t}}\right)\leq\frac{1}{2}\nu(V)$ we terminate with case (2b), returning the cut $\left(V\setminus\bigcup_{t=0}^{T_{0}}{S_{t}},\bigcup_{t=0}^{T_{0}}{S_{t}}\right)$.
Next, assume that $\nu\left(\bigcup_{t=0}^{T_{0}}{S_{t}}\right)>\frac{1}{2}\nu(V)$. This means that $\nu\left(S_{T_{0}}\right)>\frac{1}{4}\nu(V)$. We set $A_{1}=V\setminus(\bigcup_{t=0}^{T_{0}-1}{S_{t}})$ and $A_{2}=A_{1}\setminus S_{T_{0}}$ and terminate in case (2c). The guarantee $\mu(A_{2})\geq\frac{1}{4}\mu(V)$ is given by Remark 3.10.
This concludes case (2) of Theorem 3.9. Next, assume Theorem 3.9 terminated with case (3). This means that with high probability $G[A]$ is an $\Omega(\phi)$-expander with respect to $\mu$. Additionally, the cut $(A,V\setminus A)$ is a result of the trimming step (see Remark 3.10) and satisfies the properties mentioned in Remark 3.10. In particular, we have $0<\mu(V\setminus A)\leq\frac{\mu(V)}{2}$. We split into cases according to $\nu(V\setminus A)$. If $\nu(V\setminus A)\geq\frac{1}{2}\nu(V)$, we terminate with case (3a). Otherwise, we terminate with case (3b).
In all cases, the required flows are given by Remark 3.10. ∎
@@
We then apply Corollary 6.5 on $G_{\tilde{S}}$ with $\phi_{S}=\Theta\left(\frac{1}{\log n\cdot f(S)}\right)$.  [^20] The reason we prefer to apply Corollary 6.5 on $G_{\tilde{S}}$ instead of $G^{\prime}[S^{\prime}]$ is that applying it on the latter may create some edge cases which complicate the argument. For example, it may turn out that an edge $e=(v,u)\in E_{S}$ has both $v$ and $u$ on the same side of the cut while $x_{e}$ is on the other. The exact constant $c_{\phi}$ in $\phi_{S}=\frac{c_{\phi}}{\log n\cdot f(S)}$ will be $c_{\phi}\leq 1$ and is defined below. We split into cases according to the result of the corollary (see Figure 6). If Corollary 6.5 terminates with case (1), we terminate.

### v2

@@ 6.1.1 Overview
Given a cluster $S\subseteq V$, we would like to check if its boundary edges $B_{S}=E(S,V\setminus S)$ expand within G[S]G[S]. [^21] The formal argument is performed in the subdivision graph (i.e. proving that $X_{B_{S}}$ expands in $G^{\prime}[S^{\prime}]$). See Section 6.1.2 for details. For this overview, we omit these details for simplicity. If they are not, we would like to find a sparse cut $(A,S\setminus A)$ with respect to the corresponding weighting $\mu_{S}=\mathbbm{1}_{B_{S}}$, that certifies that the boundary edges do not expand. We then recur on both $A$ and $S\setminus A$ (with the new weightings $\mu_{A}=\mathbbm{1}_{B_{A}}$ and similarly $\mu_{S\setminus A}=\mathbbm{1}_{B_{S\setminus A}}$). We stop when we find a sub-cluster in which the boundary edges expand (equivalently, the sub-cluster expands with respect to the weighting given by its boundary edges). This can be facilitated using Theorem 3.9. Note that when we recur into a sub-cluster (e.g., $A\subseteq S$), the new cut edges $E(S,A\setminus S)$ are now considered boundary edges of the sub-clusters $A$ and $S\setminus A$. However, because the cut was sparse, the capacity of these cut edges is small compared to the boundary capacity already present in $S$. Consequently, for each sub-cluster, the total capacity of its boundary edges decreases relative to that of its parent cluster, leading to a polylogarithmic recursion depth. Note that we must carefully adjust the target expansion $\phi$ when we invoke Theorem 3.9 on the different sub-clusters, to ensure that the final sub-clusters $S_{i}\subseteq S$, $\Omega\!\left(\frac{1}{f(S_{i})\cdot\log n}\right)$-expand with respect to their boundary edges.
@@
We can imagine the sequence of cuts and recursions as a binary tree whose root is $S$, and where in each step, the children of the current node are the two sides of the cut. Because the cut is balanced (by Theorem 3.9), we can ensure that the depth of the tree is poly-logarithmic, so the total running time is $\tilde{O}(m)$.
@@
As for the second result of Theorem 6.2, we would like to use the flows that are promised by Theorem 3.10 to route demands from the cut edges to the boundary edges. Specifically, in each iteration we route demands from the cut edges $E(A,D\setminus A)$, where $D\subseteq S$ denotes the current sub-cluster of $S$, to the boundary edges of $A$ or of $D\setminus A$. In this manner, we could route the flow from the inter-cluster edges $F$ to the boundary edges $B_{S}$ by going over the binary tree from the leaves upwards: each time we encounter a cut $(A,D\setminus A)$ of the current sub-cluster $D\subseteq S$, we route the mass currently residing on the cut edges to the boundary edges of either $A$ or $D\setminus A$. Note that the demands on the boundary edges accumulate as we perform this step. In particular, after accounting for all relevant details, we obtain that routing a unit of flow from each edge of $E(A,D\setminus A)$ to $A$ adds $O\left(\frac{1}{f(D)}\right)$ units [^22] Recall that $f(D)\mathrel{\mathop{:}}=c_{f}\log\!\left(\frac{\sigma}{|D|}\right)\!\cdot\log\log n$. of load to the boundary edges of $A$ (and similarly if we route to the boundary of $D\setminus A$). Assuming the load on the boundary edges of $A$ and on the cut edges $E(A,D\setminus A)$ (accumulated in the lower levels of the binary tree) was $\alpha$, this means that when we go upwards in the binary tree, the load on the boundary of $D$ will now be $\alpha^{\prime}=\alpha\cdot\left(1+\frac{O(1)}{f(D)}\right)$.
@@
If we could always route to the side with the smaller number of vertices (i.e., route to $A$ if $|A|\leq|D\setminus A|$ and to $|D\setminus A|$ otherwise), this approach would work. The reasoning is as follows. Since $f(D)=c_{f}\cdot\log\left(\frac{\sigma}{|D|}\right)\log\log n$ and $\sigma\geq\frac{3}{2}|S|$, it follows that if $|D|\leq|S|\cdot 2^{-j}$ for some $j\geq 1$, then $f(D)=\Omega(j\log\log n)$. This ensures that for each edge $e$, if we consider the sequence of clusters $D_{i}$ (along the binary tree) in which the load on $e$ increases (i.e., when $e$ is a boundary edge on the side with fewer vertices), the product $\prod_{i}{\left(1+\frac{1}{f(D_{i})}\right)}$ behaves like $\prod_{j=1}^{\log n}{\left(1+\frac{1}{j\log\log n}\right)}$, so it remains bounded by a constant. Hence, the total load accumulated on each edge, determined by this product, is constant (see Claim A.3).
@@
However, there is a problem with this approach. Specifically, as detailed in Theorem 3.10, due to the algorithms implying Theorem 3.9, the returned sparse cut $(A,D\setminus A)$ of a sub-cluster $D\subseteq S$ is created from a sequence of cuts $\{S_{t}\}_{t=0}^{T-1}$, where if we define $A_{0}=D$, $A_{t+1}=A_{t}\setminus S_{t}$ ($t\in\{0,\ldots,T-1\}$), we have that $S_{t}\subseteq A_{t}$ is the cut we remove from the leftover cluster at step $t$. All of these cuts are sparse (i.e., $(A_{t}\setminus S_{t},S_{t})$ is a sparse cut of $A_{t}$), and additionally we can route from the cut edges to both sides ($S_{t}$ and $A_{t}\setminus S_{t}$). Ultimately, when $A=D\setminus\left(\bigcup_{t}{S_{t}}\right)$, we can route flow from the cut edges $(A,D\setminus A)$ to the side $D\setminus A$ (see Figure 6(2a)(2a)). However, if we attempt to route instead towards $A$, we may incur congestion on the same edges multiple times (specifically, $T$ times), [^23] This happens since $\{S_{t}\}_{t=0}^{T-1}$ are disjoint while $\{A_{t}\setminus S_{t}\}_{t=0}^{T-1}$ are not. Therefore the routings from the cuts $(A_{t}\setminus S_{t},S_{t})$ to $A_{t}\setminus S_{t}$ may overlap and accumulate congestion. preventing a routing with sufficiently low congestion. We emphasize that this is not an issue in $D\setminus A$, because each cut is removed from the graph after it is processed: we route from $E(A_{t}\setminus S_{t},S_{t})$ to $S_{t}$ within $G[S_{t}]$, and then $S_{t}$ is removed from the remaining graph. As a result, the corresponding flow routings remain disjoint (see Figure 6(2a)(2a)) and the congestion does not accumulate.
@@
This routing is good for us if $|D\setminus A|\leq|A|$, but we have no such guarantee from Theorem 3.9 (as the theorem only bounds $\mu(D\setminus A)$, which is the number of boundary edges). Instead, we choose to do something slightly more nuanced: we stop the sequence of cuts $\{S_{t}\}_{t=0}^{T-1}$ early at some $T_{0}<T$, as soon as its size $\left|\bigcup_{t=0}^{T_{0}}{S_{t}}\right|$ is a constant fraction of $|D|$. In other words, we find the first $T_{0}$ such that $\left|\bigcup_{t=0}^{T_{0}}{S_{t}}\right|=\Theta(|D|)$. Denote $A=D\setminus\left(\bigcup_{t=0}^{T_{0}}{S_{t}}\right)$. Assuming that at this step, both sides of the cut are of size $\Theta(|D|)$, this is enough to ensure that the recursion depth is poly-logarithmic (even though $\mu(D\setminus A)=\mu\left(\bigcup_{t=0}^{T_{0}}{S_{t}}\right)$ may be arbitrarily small), and allows us to bound $|D\setminus A|=\left|\bigcup_{t=0}^{T_{0}}{S_{t}}\right|=\Theta(|D|)$ (see Figure 6(2b)(2b)).
@@
The final obstacle with this approach is that it may be that while $\left|\bigcup_{t=0}^{T_{0}-1}{S_{t}}\right|$ is small (say $o(|D|)$), it’s possible $\left|\bigcup_{t=0}^{T_{0}}{S_{t}}\right|$ is already too large (say $(1-o(1))\cdot|D|$), so $\left|\bigcup_{t=0}^{T_{1}}{S_{t}}\right|$ is not $\Theta(|D|)$ for any $T_{1}$. In this case we perform two steps: we denote $A_{1}\mathrel{\mathop{:}}=D\setminus\left(\bigcup_{t=0}^{T_{0}-1}{S_{t}}\right)$, which is a cut of $D$, and then further cut $A_{1}$ into $(S_{T_{0}},A_{2}\mathrel{\mathop{:}}=A_{1}\setminus S_{T_{0}})$. Routing from $E(A_{1},D\setminus A_{1})$ to boundary of the side $D\setminus A_{1}$ is done like before (as mentioned, routing to the side $D\setminus A_{1}$ does not incur multiple congestion on the same edges, because in $D\setminus A_{1}=\bigcup_{t=0}^{T_{0}-1}{S_{t}}$, the cuts are removed at each step). Routing from $E(D\setminus A_{2},A_{2})$ to the boundary of $A_{2}$ is possible because $S_{T_{0}}$ now represents a single internal cut rather than a sequence of cuts. Consequently, we do not incur the congestion multiple times (see Figure 6(2c)(2c)).
@@
In total, this partitioning scheme guarantees that we always route from the cut edges to the boundary edges of a side whose size is at most $c\cdot|D|$. This suffices to show that the total accumulated load on each boundary edge remains bounded by a constant.
@@
We are going to apply Theorem 3.9 on the cluster $S\subseteq V$. However, as discussed in Section 6.1.1, Theorem 3.9 accounts only for the vertex measure $\mu$. In our setting, we also need to control the number of vertices in the relevant sets, and therefore require a refined variant that incorporates both $\mu$ and cardinality. The following corollary of Theorems 3.9 and 3.10 provides this. For generality, we state it with an arbitrary measure $\nu:V\rightarrow\mathop{\mathbb{R}}_{\geq 0}$, even though in our application $\nu$ will always be the counting measure $\nu(A)=|A|$. [^24] As will be described below, we use Corollary 6.5 on the subdivision graph and $\nu(A)$ will actually be the number of “non-split vertices” in $A$, i.e., $\nu(A)=|A\cap V|$. Recall the notion of routing from a cut to a set from Section 3.3. See Figure 6 for an illustration of the cases of the corollary.
@@
Given a graph $G=(V,E)$ of $m$ edges, a parameter $\phi>0$, and two weight functions $\mu,\nu:V\to\mathop{\mathbb{R}}_{\geq 0}$ on the vertices, there exists a randomized algorithm which takes $\tilde{O}(m)$ time and must end in one of the following cases:
@@
    We find a cut $(A,V\setminus A)$ of $V$ with $\mu$-expansion $\Phi^{\mu}_{G}(A,V\setminus A)=O(\phi\log n)$, $\mu(V\setminus A)=\Omega\left(\frac{\mu(V)}{\log n}\right)$, and $\nu(V\setminus A)\leq\frac{1}{4}\nu(V)$. Additionally, there exists a flow routable with congestion $O(1)$ in $G[V\setminus A]$, such that each cut edge of $E(A,V\setminus A)$ is a source of one unit of flow [^25] Formally, there is a flow in $G[V\setminus A]$ routable with congestion $O(1)$, such that each $v\in V\setminus A$ sends $\deg_{G}(v)-\deg_{G[V\setminus A]}(v)$ units of flow and receives at most $O(\phi\log n\cdot\mu(v))$ units. The rest of the flows in this corollary are defined similarly. and each $v\in V\setminus A$ receives at most $O(\phi\log n\cdot\mu(v))$ units.
@@
    We find a cut $(A,V\setminus A)$ of $V$ with $\mu$-expansion $\Phi^{\mu}_{G}(A,V\setminus A)=O(\phi\log n)$, and $\frac{1}{4}\nu(V)\leq\nu(V\setminus A)\leq\frac{1}{2}\nu(V)$. Additionally, there exists a flow routable with congestion $O(1)$ in $G[V\setminus A]$, such that each cut edge of $E(A,V\setminus A)$ is a source of one unit of flow and each $v\in V\setminus A$ receives at most $O(\phi\log n\cdot\mu(v))$ units.
@@
    We find a cut $(A_{1},V\setminus A_{1})$ of $V$ and a cut $(A_{2},A_{1}\setminus A_{2})$ of $A_{1}$, with $\mu$-expansion $\Phi^{\mu}_{G}(A_{1},V\setminus A_{1}),\Phi^{\mu}_{G}(A_{2},A_{1}\setminus A_{2})=O(\phi\log n)$. Additionally, $\mu(A_{2})\geq\frac{1}{4}\mu(V)$ and $\nu(A_{1}\setminus A_{2})\geq\frac{1}{4}\nu(V)$. [^26] In particular, the sets $V\setminus A_{1},A_{2},$ and $A_{1}\setminus A_{2}$ are all small with respect to either $\mu$ or $\nu$. There exists a flow routable with congestion $O(1)$ in $G[V\setminus A_{1}]$, such that each cut edge of $E(A_{1},V\setminus A_{1})$ is a source of one unit of flow and each $v\in V\setminus A_{1}$ receives at most $O(\phi\log n\cdot\mu(v))$ units. There also exists a flow routable with congestion $O(1)$ in $G[A_{2}]$, such that each cut edge of $E(A_{2},A_{1}\setminus A_{2})$ is a source of one unit of flow and each $v\in A_{2}$ receives at most $O(\phi\log n\cdot\mu(v))$ units.
@@
    We find a cut $(A,V\setminus A)$ with $\mu$-expansion $\Phi^{\mu}_{G}(A,V\setminus A)=O(\phi\log n)$, $0<\mu(V\setminus A)\leq\frac{\mu(V)}{2}$, and $\nu(V\setminus A)\geq\frac{1}{2}\nu(V)$. Additionally, there exists a flow routable with congestion $O(1)$ in $G[A]$, such that each cut edge of $E(A,V\setminus A)$ is a source of one unit of flow and each $v\in A$ receives at most $O(\phi\cdot\mu(v))$ units.
@@
    We find a cut $(A,V\setminus A)$ with $\mu$-expansion $\Phi^{\mu}_{G}(A,V\setminus A)=O(\phi\log n)$, with high probability, $G[A]$ is an $\Omega(\phi)$-expander with respect to $\mu$, and $\nu(V\setminus A)\leq\frac{1}{2}\nu(V)$. Additionally, there exists a flow routable with congestion $O(1)$ in $G[V\setminus A]$, such that each cut edge of $E(A,V\setminus A)$ is a source of one unit of flow and each $v\in V\setminus A$ receives at most $O(\phi\log n\cdot\mu(v))$ units. Furthermore, there exists a flow routable with congestion $O(1)$ in $G[A]$, such that each cut edge of $E(A,V\setminus A)$ is a source of one unit of flow and each $v\in A$ receives at most $O(\phi\cdot\mu(v))$ units.
@@
We apply Theorem 3.9 on $G$. If it terminated with Case (1), we are done, since this exactly matches Case (1) of the statement. Otherwise, first assume it terminated with Case (2). Recall from Theorem 3.10 that $\bar{A}$ is given as $\bar{A}=\bigcup_{t=0}^{T-1}{S_{t}}$, where if we look at the sequence of vertex sets $A_{0}=V,A_{t+1}=A_{t}\setminus S_{t}$, each $(A_{t+1},S_{t})$ is a sparse cut. We split into cases according to $\nu(\bar{A})$. If $\nu(\bar{A})\leq\frac{1}{4}\nu(V)$, we terminate with case (2a) (see Figure 6).
So assume that $\nu(\bar{A})>\frac{1}{4}\nu(V)$. We consider the first $0\leq T_{0}<T$ such that $\nu\left(\bigcup_{t=0}^{T_{0}}{S_{t}}\right)>\frac{1}{4}\nu(V)$ (i.e., $\nu\left(\bigcup_{t=0}^{T_{0}-1}{S_{t}}\right)\leq\frac{1}{4}\nu(V)$). If $\nu\left(\bigcup_{t=0}^{T_{0}}{S_{t}}\right)\leq\frac{1}{2}\nu(V)$ we terminate with case (2b), returning the cut $\left(V\setminus\bigcup_{t=0}^{T_{0}}{S_{t}},\bigcup_{t=0}^{T_{0}}{S_{t}}\right)$.
Next, assume that $\nu\left(\bigcup_{t=0}^{T_{0}}{S_{t}}\right)>\frac{1}{2}\nu(V)$. This means that $\nu\left(S_{T_{0}}\right)>\frac{1}{4}\nu(V)$. We set $A_{1}=V\setminus(\bigcup_{t=0}^{T_{0}-1}{S_{t}})$ and $A_{2}=A_{1}\setminus S_{T_{0}}$ and terminate in case (2c). The guarantee $\mu(A_{2})\geq\frac{1}{4}\mu(V)$ is given at the beginning of Theorem 3.10.
This concludes case (2) of Theorem 3.9.
Now assume that Theorem 3.9 terminated with Case (3). This means that with high probability $G[A]$ is an $\Omega(\phi)$-expander with respect to $\mu$. Additionally, the cut $(A,V\setminus A)$ is a result of the trimming step (see Theorem 3.10) and satisfies the properties mentioned in Theorem 3.10. In particular, we have $0<\mu(V\setminus A)\leq\frac{\mu(V)}{2}$. We split into cases according to $\nu(V\setminus A)$. If $\nu(V\setminus A)\geq\frac{1}{2}\nu(V)$, we terminate with case (3a). Otherwise, we terminate with case (3b).
In all cases, the required flows are given directly by Theorem 3.10. ∎
@@
We then apply Corollary 6.5 on $G_{\tilde{S}}$ with $\phi_{S}=\Theta\left(\frac{1}{\log n\cdot f(S)}\right)$. [^27] The reason we prefer to apply Corollary 6.5 on $G_{\tilde{S}}$ instead of $G^{\prime}[S^{\prime}]$ is that applying it on the latter may create some edge-cases which complicate the argument. For example, it may turn out that an edge $e=(v,u)\in E_{S}$ has both $v$ and $u$ on the same side of the cut while $x_{e}$ is on the other. The exact constant $c_{\phi}$ in $\phi_{S}=\frac{c_{\phi}}{\log n\cdot f(S)}$ will be $c_{\phi}=\min\{1,\frac{1}{16c_{\star}}\}$, where $c_{\star}$ is the constant governing the sparsity bounds in Corollary 6.5 (see below for details). We split into cases according to the result of the corollary (see Figure 6). If Corollary 6.5 terminates with Case (1), we terminate.